what does x and y =
the equation is x-4y=8
the other one is x=-1+y
: what does x and y =
the equation is x-4y=8
the other one is x=-1+y

Answer by BrittanyM(80)

We are given:

x - 4y = 8

and

x = -1 + y

Since we know what the value of x is in terms of y, we may solve for the values by using substitution.


( -1 + y ) - 4y = 8

-1 -3y = 8

-3y = 9

y = -3

Now that we have the value of y, we can just plug it in to find x

x - 4( -3 ) = 8

x + 12 = 8

x = -4